# thanks: Corresponding author

### v1

@@ V Summary
In this work, we use the QMETTS approach to study the thermal properties of a massive Thirring model, which is a soluble model for studying the interactions of fermions. The observables in interest are the chiral condensation which is related to the chiral phase transition, and the fermion number which corresponds to the topological charge in a sine-Gordon model dual to the massive Thirring model.
@@
The results of the chiral condensation demonstrate that the chiral symmetry is broken at high temperatures, or at low temperatures and smaller coupling constants for the massive Thirring model. At low temperatures and larger coupling constants, the chiral symmetry is restored. This indicates the existence of a chiral phase transition in the massive Thirring model. Furthermore, it can be observed that at low temperatures, the fermion number of the massive Thirring model is an integer and exhibits quantized variation with the interaction. This suggests that the ground state of the corresponding sine-Gordon model is situated within a distinct topological sector, indicating a topological phase transition. Additionally, a quantized alteration in chiral condensation with fermion number is also observed at low temperatures.
@@
A comparison of the results obtained from the QMETTS and exact diagonalization demonstrates that the utilisation of QMETTS is capable of reproducing the outcomes of the exact diagonalization. In particular, with regard to the low-temperature case, it is evident that the QMETTS error is not accumulated in line with the evolution of $\beta$. This indicates that the QMETTS remains a valid tool for the study of low-temperature systems, as well as physical systems exhibiting topological phase transitions. In light of these findings, this work offers a valuable reference point for future QMETTS or QITE-based studies. As quantum computing technology advances, it is plausible that one will be able to surmount the sign problem in lattice QCD with the aid of quantum computing.
@@ Appendix A Details on the $\mathcal{O}(a)$ terms
$\begin{split}&\partial_{h}\chi(n)=\frac{1}{a}\left(\chi(n+2)-\chi(n)\right),\\ \end{split}$ (28)
@@
$\begin{split}&\sum_{h}\bar{\psi}(h)i\gamma_{1}\partial_{x}\psi(h)\\ &=-\frac{i}{a^{2}}\sum_{n}\left(\chi^{\dagger}(n)\chi(n+1)-\chi^{\dagger}(n+1)\chi(n)\right)\\ \end{split}$ (29)
$\begin{split}&a\sum_{h}\bar{\psi}(h)\psi(h)=\sum_{n}(-1)^{n}\chi^{\dagger}(n)\chi(n)\\ \end{split}$ (30)
$\begin{split}&a\sum_{h}\bar{\psi}(h)\gamma_{0}\psi(h)=\sum_{n}\chi^{\dagger}(n)\chi(n)\\ \end{split}$ (31)
$\begin{split}&a^{2}\sum_{h}(\bar{\psi}(h)\psi(h))^{2}=\\ &\sum_{n}\left\{(\chi^{\dagger}(n)\chi(n))^{2}-(\chi^{\dagger}(n)\chi(n))(\chi^{\dagger}(n+1)\chi(n+1))\right\}\\ &-a\sum_{even\;n}\left\{(\partial_{h}\chi^{\dagger}(n))\chi(n)\chi^{\dagger}(n+1)\chi(n+1)\right.\\ &\left.+\chi^{\dagger}(n+2)(\partial_{h}\chi(n))\chi^{\dagger}(n+1)\chi(n+1)\right\}\\ \end{split}$ (32)
$\begin{split}&a^{2}\sum_{h}(\bar{\psi}(h)\gamma_{0}\psi(h))^{2}=\\ &\sum_{n}\left((\chi^{\dagger}(n)\chi(n))^{2}+(\chi^{\dagger}(n)\chi(n))(\chi^{\dagger}(n+1)\chi(n+1))\right)\\ &+a\sum_{even\;n}\left\{(\partial_{h}\chi^{\dagger}(n))\chi(n)\chi^{\dagger}(n+1)\chi(n+1)\right.\\ &\left.+\chi^{\dagger}(n+2)(\partial_{h}\chi(n))\chi^{\dagger}(n+1)\chi(n+1)\right\}\\ \end{split}$ (33)
$\begin{split}&a^{2}\sum_{h}(\bar{\psi}(h)\gamma_{1}\psi(h))^{2}=\\ &\frac{1}{2}\sum_{n}\left(\chi^{\dagger}(n)\chi(n+1)+\chi^{\dagger}(n+1)\chi(n)\right)^{2}\\ &-\frac{a}{2}\sum_{even\;n}\left((\partial_{h}\chi^{\dagger}(n))\chi(n+1)\chi^{\dagger}(n)\chi(n+1)\right.\\ &\left.+\chi^{\dagger}(n+2)\chi(n+1)(\partial_{h}\chi^{\dagger}(n))\chi(n+1)\right.\\ &\left.+\chi^{\dagger}(n+1)(\partial_{h}\chi(n))\chi^{\dagger}(n)\chi(n+1)\right.\\ &\left.+\chi^{\dagger}(n+1)\chi(n+2)\chi^{\dagger}(n+1)(\partial_{h}\chi(n))\right)\\ &-\frac{a}{2}\sum_{even\;n}\left\{(\partial_{h}\chi^{\dagger}(n))\chi(n+1)\chi^{\dagger}(n)\chi(n+1)\right.\\ &\left.+\chi^{\dagger}(n+1)\chi(n+2)\chi^{\dagger}(n+1)(\partial_{h}\chi(n))\right.\\ &\left.+\chi^{\dagger}(n+1)(\partial_{h}\chi(n))\chi^{\dagger}(n)\chi(n+1)\right.\\ &\left.+\chi^{\dagger}(n+1)\chi(n+2)(\partial_{h}\chi^{\dagger}(n))\chi(n+1)\right\}\\ \end{split}$ (34)
Ignoring the $\mathcal{O}(a)$ terms,
$\begin{split}&a^{2}\sum_{h}(\bar{\psi}(h)\psi(h))^{2}=\\ &\sum_{n}(\chi^{\dagger}(n)\chi(n)\left(\chi^{\dagger}(n)\chi(n)-\chi^{\dagger}(n+1)\chi(n+1)\right)+\mathcal{O}(a)\\ \end{split}$ (35)
$\begin{split}&a^{2}\sum_{h}(\bar{\psi}(h)\gamma_{0}\psi(h))^{2}=\\ &\sum_{n}(\chi^{\dagger}(n)\chi(n)\left(\chi^{\dagger}(n)\chi(n)+\chi^{\dagger}(n+1)\chi(n+1)\right)+\mathcal{O}(a)\\ \end{split}$ (36)
$\begin{split}&a^{2}\sum_{h}(\bar{\psi}(h)\gamma_{1}\psi(h))^{2}=\\ &\frac{1}{2}\sum_{n}\left(\chi^{\dagger}(n)\chi(n+1)+\chi^{\dagger}(n+1)\chi(n)\right)^{2}+\mathcal{O}(a)\\ \end{split}$ (37)
Then Eqs. (7-12) can be obtained by substituting $\chi$ with Eq. (6).
@@ Appendix B Measurements to obtain $S$ and $b$
Since $|\phi(\beta)\rangle$ is a quantum state, the matrix elements of $S$ and $b$ can not be obtained classically. Therefore, the matrix elements of $S$ and $b$ should be obtained by measurement. Since the Hamiltonian can also been expanded using $H=\sum_{j}h_{j}\hat{\sigma}_{j}$, essentially, one needs a quick construction of $\hat{\sigma}_{j_{1}}\hat{\sigma}_{j_{2}}+\left(\hat{\sigma}_{j_{1}}\hat{\sigma}_{j_{2}}\right)^{T}$ and $-{\rm i}\left(\hat{\sigma}_{j_{1}}\hat{\sigma}_{j_{2}}-\hat{\sigma}_{j_{2}}\hat{\sigma}_{j_{1}}\right)$. Defining,
$\begin{split}M_{a}=\begin{pmatrix}0&1&2&3\\ 1&0&3&2\\ 2&3&0&1\\ 3&2&1&0\end{pmatrix},\;\;&M_{b}=\begin{pmatrix}0&0&0&0\\ 0&0&1&1\\ 0&1&0&1\\ 0&1&1&0\end{pmatrix},\\ M_{c}=\begin{pmatrix}0&0&0&0\\ 0&0&1&0\\ 0&0&0&1\\ 0&1&0&0\end{pmatrix},\;\;&M_{d}=\begin{pmatrix}0&0&1&0\\ 0&0&0&1\\ 1&0&0&0\\ 0&1&0&0\end{pmatrix}.\\ \end{split}$ (38)
@@
$\begin{split}&\hat{\sigma}_{j_{1}}\hat{\sigma}_{j_{2}}+\left(\hat{\sigma}_{j_{1}}\hat{\sigma}_{j_{2}}\right)^{T}=(-1)^{\sum_{i_{1}}\left(M_{c}^{T}\right)^{x_{i_{1}},y_{i_{1}}}}{\rm i}^{\sum_{i_{2}}M_{b}^{x_{i_{2}},y_{i_{2}}}}\\ &\times\left(1+(-1)^{\sum_{i_{3}}M_{d}^{x_{i_{3}},y_{i_{3}}}}\right)\prod_{l}\sigma^{M_{a}^{x_{l},y_{l}}}(l),\end{split}$ (39)
@@
$\begin{split}&-{\rm i}\left(\hat{\sigma}_{j_{1}}\hat{\sigma}_{j_{2}}-\hat{\sigma}_{j_{2}}\hat{\sigma}_{j_{1}}\right)=(-1)^{\sum_{i_{1}}M_{c}^{x_{i_{1}},y_{i_{1}}}}{\rm i}^{\left(1+\sum_{i_{2}}M_{b}^{x_{i_{2}},y_{i_{2}}}\right)}\\ &\times\left(1-(-1)^{\sum_{i_{3}}M_{b}^{x_{i_{3}},y_{i_{3}}}}\right)\prod_{l}\sigma^{M_{a}^{x_{l},y_{l}}}(l),\end{split}$ (40)
where $M^{x,y}$ is the matrix element of $M$ at index $(x,y)$. $S+S^{T}$ in Eq. (18) can be obtained directly using Eq. (39). For the components of $b$ in Eq. (19), $b_{j}=\sum_{j^{\prime}}h_{j^{\prime}}\left\langle\phi|-{\rm i}\left(\hat{\sigma}_{j}\hat{\sigma}_{j^{\prime}}-\hat{\sigma}_{j^{\prime}}\hat{\sigma}_{j}\right)|\phi\right\rangle/\sqrt{C(\Delta\beta)}$, where $h_{j^{\prime}}$ are the coefficients in the expansion of Hamiltonian.

### v2

@@ 6 Summary
In this work, we use the QMETTS approach to study the thermal properties of a massive Thirring model, which is a solvable system describing fermion interactions. The key observables are the chiral condensate, which is associated with the chiral phase transition, and the fermion number, which corresponds to the topological charge in the dual sine-Gordon model.
@@
The numerical results are obtained by using qiskit to simulate the circuit on a classical computer, and with key components such as the operator coefficients and normalization factors computed exactly rather than through quantum measurements. The results of the chiral condensation demonstrate that the chiral symmetry is broken at high temperatures, or at low temperatures and smaller coupling constants for the massive Thirring model. At low temperatures and large coupling constants, the chiral symmetry is restored. These results confirm the existence of a chiral phase transition in the massive Thirring model. Furthermore, at low temperatures, the fermion number assumes integer values and exhibits quantized variation with the interaction strength. This suggests that the ground state of the corresponding sine-Gordon model is situated within a distinct topological sector, indicating a topological phase transition. Additionally, a quantized alteration in the chiral condensation along with the fermion number is also observed at low temperatures.
@@
A comparison of the results obtained from the QMETTS and exact diagonalization demonstrates that the utilization of QMETTS is capable of reproducing the outcomes of the exact diagonalization. In particular, with regard to the low-temperature case, it is evident that the QMETTS error do not accumulate with the evolution of $\beta$. This indicates that the QMETTS remains a valid tool for the study of low-temperature systems, as well as physical systems exhibiting topological phase transitions. In light of these findings, this work offers a valuable reference point for future QMETTS or QITE-based studies. As quantum computing technology advances, it is plausible that one will be able to surmount the sign problem in lattice QCD with the aid of quantum computing.
@@ Appendix A Details on the $\mathcal{O}(a)$ terms
[…]
@@
[…]
Neglecting the $\mathcal{O}(a)$ terms,
[…]
Then Eqs. (7) and (8) can be obtained by substituting $\chi$ with Eq. (6).
@@ Appendix B Measurements to obtain $S$ and $b$
Since $|\phi(\beta)\rangle$ is a quantum state, the matrix elements of $S$ and $b$ cannot be obtained classically. Therefore, the matrix elements of $S$ and $b$ should be determined through measurements. Since the Hamiltonian can also be expanded using $aH=\sum_{j}h_{j}\hat{\sigma}_{j}$, essentially, one needs a quick construction of $\hat{\sigma}_{j_{1}}\hat{\sigma}_{j_{2}}+\left(\hat{\sigma}_{j_{1}}\hat{\sigma}_{j_{2}}\right)^{T}$ and ${\rm i}\left(\hat{\sigma}_{j_{1}}\hat{\sigma}_{j_{2}}-\hat{\sigma}_{j_{2}}\hat{\sigma}_{j_{1}}\right)$. Defining,
[…]
@@
$\begin{split}&\hat{\sigma}_{j_{1}}\hat{\sigma}_{j_{2}}+\left(\hat{\sigma}_{j_{1}}\hat{\sigma}_{j_{2}}\right)^{T}=(-1)^{\sum_{i_{1}}\left(M_{c}^{T}\right)^{x_{i_{1}},y_{i_{1}}}}{\rm i}^{\sum_{i_{2}}M_{b}^{x_{i_{2}},y_{i_{2}}}}\left(1+(-1)^{\sum_{i_{3}}M_{d}^{x_{i_{3}},y_{i_{3}}}}\right)\prod_{l}\sigma^{M_{a}^{x_{l},y_{l}}}(l),\end{split}$ (29)
@@
$\begin{split}&-{\rm i}\left(\hat{\sigma}_{j_{1}}\hat{\sigma}_{j_{2}}-\hat{\sigma}_{j_{2}}\hat{\sigma}_{j_{1}}\right)=(-1)^{\sum_{i_{1}}M_{c}^{x_{i_{1}},y_{i_{1}}}}{\rm i}^{\left(1+\sum_{i_{2}}M_{b}^{x_{i_{2}},y_{i_{2}}}\right)}\left(1-(-1)^{\sum_{i_{3}}M_{b}^{x_{i_{3}},y_{i_{3}}}}\right)\prod_{l}\sigma^{M_{a}^{x_{l},y_{l}}}(l),\end{split}$ (30)
where $M^{x,y}$ is the matrix element of $M$ at index $(x,y)$. $S+S^{T}$ in Eq. (14) can be obtained directly using Eq. (29). For the components of $b$ defined in Eq. (15), $b_{j}=\sum_{j^{\prime}}h_{j^{\prime}}\left\langle\phi|-{\rm i}\left(\hat{\sigma}_{j}\hat{\sigma}_{j^{\prime}}-\hat{\sigma}_{j^{\prime}}\hat{\sigma}_{j}\right)|\phi\right\rangle/\sqrt{C(\Delta\beta)}$, where $h_{j^{\prime}}$ are the dimensionless coefficients in the expansion of Hamiltonian.